# Probing $\alpha$ clustering in $^{12}\mathrm{C}$ at CSR energies using the Jet AA Microscopic Transport Model

### v1

@@ Abstract
We investigate the sensitivity of low-energy nuclear collisions to intrinsic nuclear structure by studying the interplay between initial-state geometry and final-state observables in C+C and C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV, relevant for experiments at the Cooling Storage Ring (CSR) facility in Lanzhou and forthcoming experiments at the High Intensity heavy-ion Accelerator Facility (HIAF) in Huizhou. Calculations are performed within the Jet AA Microscopic Transport Model (JAM) using Woods–Saxon and triangular $\alpha$-clustered configurations for the $^{12}\mathrm{C}$ nucleus. The initial geometry is characterized in terms of transverse size, compactness, eccentricities, and their event-by-event fluctuations. We find that $\alpha$ clustering leads to a more compact participant configuration than the Woods–Saxon case, while transverse-size and eccentricity fluctuations show only weak sensitivity to clustering. At this beam energy, radial observables remain sensitive to geometric compactness, with the proton mean transverse momentum $\langle p_{T}\rangle$ enhanced for $\alpha$-clustered configurations, whereas pions show little sensitivity. The anisotropic response is examined using flow harmonic coefficients. We find an enhancement of the mean flow magnitudes, $\langle v_{n}\rangle=\sqrt{\langle v_{n}^{2}\rangle}$, for $\alpha$-clustered configurations at large $N_{\mathrm{part}}$, while the event-by-event fluctuation strength of individual harmonics remains small. Symmetric cumulants of the initial-state eccentricities show sensitivity to clustering, whereas the corresponding correlations among final-state flow harmonics do not exhibit a comparably strong separation. These results indicate that radial observables and correlation-based flow measurements provide complementary probes of $\alpha$ clustering in low-energy nuclear collisions.
@@ II.1 Clustering implementation
To investigate the role of nuclear structure effects, two distinct initial configurations are employed for the $^{12}\mathrm{C}$ nucleus. In the reference configuration, nucleon positions are sampled from a spherical Woods–Saxon distribution. In the alternative configuration, $^{12}\mathrm{C}$ is modeled as an $\alpha$-clustered system consisting of three $\alpha$ clusters arranged at the vertices of an equilateral triangle in the transverse plane. This triangular configuration is motivated by nuclear structure studies, which have long predicted a pronounced three-$\alpha$ cluster structure in the ground state of $^{12}\mathrm{C}$ Hafstad and Teller (1938); Harrington (1966); Ikeda et al. (1968); Noble (1970); Bertsch and Bertozzi (1971). Both transport and hydrodynamic models have been extensively employed to explore possible experimental signatures of nuclear clustering in heavy-ion collisions over a wide range of beam energies Broniowski and Ruiz Arriola (2014); He et al. (2014); Rybczyński et al. (2018); Zhang et al. (2017); Guo et al. (2017); Yao et al. (2025); Ma et al. (2020).
@@
Figure 2: The transverse size ($\langle r^{2}\rangle$), compactness ($C$) and event-by-event transverse-size fluctuations as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+C collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
Figure 3: The transverse size ($\langle r^{2}\rangle$), compactness ($C$) and event-by-event transverse-size fluctuations as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV
@@ III.1 Correlation in initial-state geometries
which provides a measure of the inverse transverse length scale governing the initial pressure gradients. Event-by-event transverse-size fluctuations are quantified by the relative width
@@
evaluated at fixed $N_{\mathrm{part}}$.
Figure 2 and  3 present scaled transverse size ($\langle r^{2}\rangle/N_{\mathrm{part}}$), compactness ($C$) and event-by-event transverse-size fluctuations as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for for C+C and C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV. In C+C collisions, the $\alpha$-clustered configuration exhibits a systematically smaller scaled transverse size than the Woods–Saxon case, indicating a more compact spatial arrangement of the participant nucleons. This reduction in $\langle r^{2}\rangle/N_{\mathrm{part}}$ is accompanied by a corresponding enhancement of the compactness over the full $N_{\mathrm{part}}$ range. In contrast, the relative transverse-size fluctuations show a weaker sensitivity to clustering, with comparable magnitudes for Woods–Saxon and $\alpha$-clustered configurations. While the magnitude of fluctuations decreases with increase in $N_{\mathrm{part}}$. The observed pattern demonstrates that, in symmetric light–light collisions, $\alpha$ clustering primarily shifts the average transverse scale of the system rather than significantly enhancing event-by-event size fluctuations.
A different behavior is observed in C+Pb collisions. While the average compactness remains larger for the $\alpha$-clustered configuration, reflecting the increased local packing of nucleons in the carbon projectile, the relative transverse-size fluctuations are comparable or slightly larger for the Woods–Saxon case. In this asymmetric system, fluctuations of the transverse size are dominated by the geometry and participant fluctuations on the Pb side, which reduces the relative impact of clustering in the lighter nucleus. Consequently, $\alpha$ clustering continues to modify the mean transverse geometry but does not lead to an enhancement of transverse-size fluctuations. Overall, the comparison shown in Figures  2 and  3 demonstrates that $\alpha$ clustering predominantly affects the average compactness of the participant zone, while its influence on transverse-size fluctuations is comparatively modest and predominant at low $N_{\mathrm{part}}$ ranges.
Figure 4: Event averaged participant eccentricities and their event-by-event fluctuations as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+C collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
@@
Event-averaged eccentricities are obtained as functions of $N_{\mathrm{part}}$ by averaging $\varepsilon_{n}$ over events within a given $N_{\mathrm{part}}$ bin. Event-by-event fluctuations are quantified using the variance
@@
Figure 4 presents the average eccentricities $\varepsilon_{n}$ and their event-by-event fluctuations as a function of $N_{\mathrm{part}}$ for C+C collisions. For both the Woods–Saxon and $\alpha$-clustered configurations, the eccentricities follow the ordering $\varepsilon_{2}>\varepsilon_{3}>\varepsilon_{1}$, with comparable magnitudes in the two cases. The relative fluctuations of each $\varepsilon_{n}$ remain approximately constant as a function of $N_{\mathrm{part}}$, with the ordering $\sigma(\varepsilon_{1})>\sigma(\varepsilon_{3})>\sigma(\varepsilon_{2})$.
Figure 5: Event averaged participant eccentricities and their event-by-event fluctuations as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV. Results are calculated using C participants only (see text for details)
Figure 6: Event averaged participant eccentricities and their event-by-event fluctuations as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
@@
Correlations between different harmonic components of the initial geometry are further investigated using mixed moments such as $\langle\varepsilon_{1}\varepsilon_{3}\rangle$ and $\langle\varepsilon_{2}\varepsilon_{3}\rangle$, as well as their squared counterparts. These observables probe the degree to which different geometric deformations are correlated on an event-by-event basis and provide information beyond that contained in individual eccentricity magnitudes.
@@
where the angular brackets denote event averages at fixed $N_{\mathrm{part}}$. Positive values of $\mathrm{SC}\{m,n\}$ indicate a positive correlation between $\varepsilon_{m}^{2}$ and $\varepsilon_{n}^{2}$, while negative values signal an anticorrelation. In this work, symmetric cumulants are evaluated for the harmonic combinations $(m,n)=(1,3)$ and $(2,3)$ as functions of $N_{\mathrm{part}}$. For presentation purposes, the cumulants are scaled by $N_{\mathrm{part}}$ to reduce the trivial dilution of correlations with increasing system size.
@@ III.3 Correlation in final-state particles
Event-by-event fluctuations of the mean transverse momentum are quantified using a two-particle correlator,
$\delta p_{T,\alpha}^{2}=\langle p_{T}p_{T}\rangle_{\alpha}-\langle p_{T}\rangle_{\alpha}^{2},$ (10)
where $\langle p_{T}p_{T}\rangle_{\alpha}$ is constructed from distinct particle pairs within the same event in order to remove self-correlations. This definition isolates dynamical fluctuations beyond trivial statistical contributions.
@@
To investigate the response to initial transverse-radius fluctuations, event-by-event dynamical transverse-momentum fluctuations are studied and quantified using a two-particle correlator,
$\delta p_{T,p}^{2}=\langle\!\langle\langle p_{T}p_{T}\rangle_{p}-\langle p_{T}\rangle_{p}^{2}\rangle\!\rangle,$ (12)
where $\langle p_{T}p_{T}\rangle_{p}=(Q_{1}^{2}-Q_{2})/[N(N-1)]$ is constructed from distinct particle pairs within the same event, with $Q_{1}=\sum_{i}p_{T,i}$ and $Q_{2}=\sum_{i}p_{T,i}^{2}$. This definition removes self-correlations and isolates dynamical fluctuations of the event-wise mean transverse momentum.
@@
Figure 16: Two particle $\langle p_{T}\rangle$ correlation of protons and pions as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+C collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
Figure 17: Two particle $\langle p_{T}\rangle$ correlation of protons and pions as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
@@
To characterize the magnitude and event-by-event fluctuations of the flow harmonics, we evaluate moments of the event-wise flow coefficients using the $Q$-vector formalism. For each event, the quantity $v_{n}^{2}=|Q_{n}|^{2}/M^{2}$ is constructed, where $Q_{n}=\sum_{i}e^{in\phi_{i}}$ and $M$ denotes the multiplicity of selected particles. The typical magnitude of the flow harmonics is characterized by $\langle v_{n}\rangle=\sqrt{\langle v_{n}^{2}\rangle}$, while the fluctuation strength of a given harmonic is quantified by the variance $\langle v_{n}^{4}\rangle-\langle v_{n}^{2}\rangle^{2}$, with angle brackets denoting an average over events at fixed $N_{\mathrm{part}}$.
@@
Figure 18: $\langle v_{1}\rangle$ and its event-by-event fluctuation as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV
@@
Figure 19: $\langle v_{2}\rangle$ and its event-by-event fluctuation as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV
@@
Figure 20: $\langle v_{3}\rangle$ and its event-by-event fluctuation as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV
@@
Figures 18, 19, and 20 present the mean flow magnitude and its event-by-event fluctuations for the first-, second-, and third-order flow harmonics. Interestingly, we observe that $\alpha$ clustering leads to a significant enhancement of the mean flow magnitude, particularly at large $N_{\mathrm{part}}$ ($\gtrsim 70$). In contrast, the event-by-event fluctuation strength of individual harmonics remains small and is not clearly distinguishable within the current statistical precision. This indicates that $\alpha$ clustering primarily modifies the average initial geometry and its collective response, while inducing only modest changes in the magnitude of single-harmonic fluctuations. Consequently, precise experimental measurements of flow observables in this regime could provide valuable constraints on the presence and strength of $\alpha$ clustering in light nuclei.
@@
Finally, symmetric cumulants of $v_{n}$ are evaluated. Figure 21 shows $\mathrm{SC}(v_{1},v_{3})$ and $\mathrm{SC}(v_{2},v_{3})$ for C+Pb collisions. Both configurations yield a positive $\mathrm{SC}\{v_{1},v_{3}\}$, with no statistically significant distinction between them. The $\mathrm{SC}(v_{2},v_{3})$ values are consistent with zero for both $\alpha$-clustering and Woods–Saxon cases.

### v2

@@ Abstract
We investigate the sensitivity of low-energy nuclear collisions to intrinsic nuclear structure by studying the interplay between initial-state geometry and final-state observables in C+C and C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV, relevant for experiments at the Cooling Storage Ring (CSR) facility in Lanzhou and forthcoming experiments at the High Intensity heavy-ion Accelerator Facility (HIAF) in Huizhou. Calculations are performed within the Jet AA Microscopic Transport Model (JAM) using Woods–Saxon and triangular $\alpha$-clustered configurations for the $^{12}\mathrm{C}$ nucleus. The initial geometry is characterized in terms of transverse size, compactness, eccentricities, and their ensemble-averaged fluctuations. We find that $\alpha$ clustering leads to a more compact participant configuration than the Woods–Saxon case, while transverse-size and eccentricity fluctuations show only weak sensitivity to clustering. At this beam energy, radial observables remain sensitive to geometric compactness, with the ensemble-averaged proton mean transverse momentum $\langle p_{T}\rangle$ enhanced for $\alpha$-clustered configurations, whereas pions show little sensitivity. The anisotropic response is examined using flow harmonic coefficients. We find an enhancement of the root-mean-square flow magnitudes, $v_{n}\{2\}=\sqrt{\langle v_{n}^{2}\rangle}$, for $\alpha$-clustered configurations at large $N_{\mathrm{part}}$, while the ensemble-averaged fluctuation strength of individual harmonics remains small. Symmetric cumulants of the initial-state eccentricities show sensitivity to clustering, whereas the corresponding ensemble-averaged correlations among final-state flow harmonics do not exhibit a comparably strong separation. These results indicate that radial observables and correlation-based flow measurements provide complementary probes of $\alpha$ clustering in low-energy nuclear collisions.
@@ II.1 Clustering implementation
To investigate the role of nuclear structure effects, two distinct initial configurations are employed for the $^{12}\mathrm{C}$ nucleus. In the reference configuration, nucleon positions are sampled from a spherical Woods–Saxon distribution. In the alternative configuration, $^{12}\mathrm{C}$ is modeled as an $\alpha$-clustered system consisting of three $\alpha$ clusters arranged at the vertices of an equilateral triangle in the transverse plane. This triangular configuration is motivated by nuclear structure studies, which have long predicted a pronounced three-$\alpha$ cluster structure in the ground state of $^{12}\mathrm{C}$ Hafstad and Teller (1938); Harrington (1966); Ikeda et al. (1968); Noble (1970); Bertsch and Bertozzi (1971). Both transport and hydrodynamic models have been extensively employed to explore possible experimental signatures of nuclear clustering in heavy-ion collisions over a wide range of beam energies Broniowski and Ruiz Arriola (2014); He et al. (2014); Rybczyński et al. (2018); Zhang et al. (2017); Guo et al. (2017); Ma et al. (2020); Liu et al. (2024); Yao et al. (2025).
@@
Figure 2: The transverse size ($\langle r^{2}\rangle$), compactness ($C$), and the corresponding transverse-size fluctuation strength as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+C collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
Figure 3: The transverse size ($\langle r^{2}\rangle$), compactness ($C$), and the corresponding transverse-size fluctuation strength as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
@@ III.1 Correlation in initial-state geometries
which provides a measure of the inverse transverse length scale governing the initial pressure gradients. Fluctuations of the transverse size are quantified by the relative width
@@
evaluated as an ensemble average at fixed $N_{\mathrm{part}}$.
Figure 2 and 3 present the scaled transverse size ($\langle r^{2}\rangle/N_{\mathrm{part}}$), compactness ($C$), and the corresponding transverse-size fluctuation strength as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+C and C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV. In C+C collisions, the $\alpha$-clustered configuration exhibits a systematically smaller scaled transverse size than the Woods–Saxon case, indicating a more compact spatial arrangement of the participant nucleons. This reduction in $\langle r^{2}\rangle/N_{\mathrm{part}}$ is accompanied by a corresponding enhancement of the compactness over the full $N_{\mathrm{part}}$ range. In contrast, the relative transverse-size fluctuations show a weaker sensitivity to clustering, with comparable magnitudes for Woods–Saxon and $\alpha$-clustered configurations, while the magnitude of the fluctuations decreases with increasing $N_{\mathrm{part}}$. The observed pattern demonstrates that, in symmetric light–light collisions, $\alpha$ clustering primarily shifts the average transverse scale of the system rather than significantly enhancing transverse-size fluctuations.
A different behavior is observed in C+Pb collisions. While the average compactness remains larger for the $\alpha$-clustered configuration, reflecting the increased local packing of nucleons in the carbon projectile, the relative transverse-size fluctuations are comparable or slightly larger for the Woods–Saxon case. In this asymmetric system, fluctuations of the transverse size are dominated by the geometry and participant fluctuations on the Pb side, which reduces the relative impact of clustering in the lighter nucleus. Consequently, $\alpha$ clustering continues to modify the mean transverse geometry but does not lead to an enhancement of transverse-size fluctuations. Overall, the comparison shown in Figs. 2 and 3 demonstrates that $\alpha$ clustering predominantly affects the average compactness of the participant zone, while its influence on transverse-size fluctuations is comparatively modest and most pronounced at low $N_{\mathrm{part}}$.
Figure 4: Ensemble-averaged participant eccentricities and their corresponding fluctuation strengths as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+C collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
@@
Event-averaged eccentricities are obtained as functions of $N_{\mathrm{part}}$ by averaging $\varepsilon_{n}$ over events within a given $N_{\mathrm{part}}$ bin. The fluctuation strength is quantified using the variance
@@
Figure 4 presents the average eccentricities $\varepsilon_{n}$ and their corresponding fluctuation strengths as a function of $N_{\mathrm{part}}$ for C+C collisions. For both the Woods–Saxon and $\alpha$-clustered configurations, the eccentricities follow the ordering $\varepsilon_{2}>\varepsilon_{3}>\varepsilon_{1}$, with comparable magnitudes in the two cases. The relative fluctuations of each $\varepsilon_{n}$ remain approximately constant as a function of $N_{\mathrm{part}}$, with the ordering $\sigma(\varepsilon_{1})>\sigma(\varepsilon_{3})>\sigma(\varepsilon_{2})$.
Figure 5: Ensemble-averaged participant eccentricities and their corresponding fluctuation strengths as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV. Results are calculated using C participants only (see text for details)
Figure 6: Ensemble-averaged participant eccentricities and their corresponding fluctuation strengths as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
@@
Correlations between different harmonic components of the initial geometry are further investigated using mixed moments such as $\langle\varepsilon_{1}\varepsilon_{3}\rangle$ and $\langle\varepsilon_{2}\varepsilon_{3}\rangle$, as well as their squared counterparts. These observables probe the degree to which different geometric deformations are correlated within an ensemble of events and provide information beyond that contained in individual eccentricity magnitudes.
@@
where the angular brackets denote ensemble averages over events at fixed $N_{\mathrm{part}}$. Positive values of $\mathrm{SC}\{m,n\}$ indicate a positive correlation between $\varepsilon_{m}^{2}$ and $\varepsilon_{n}^{2}$, while negative values signal an anticorrelation. In this work, symmetric cumulants are evaluated for the harmonic combinations $(m,n)=(1,3)$ and $(2,3)$ as functions of $N_{\mathrm{part}}$. For presentation purposes, the cumulants are scaled by $N_{\mathrm{part}}$ to reduce the trivial dilution of correlations with increasing system size.
@@ III.3 Correlation in final-state particles
Transverse-momentum fluctuations are quantified using the variance of the single-particle transverse-momentum distribution within a fixed $N_{\mathrm{part}}$ class,
$\delta p_{T,\alpha}^{2}=\langle p_{T}^{2}\rangle_{\alpha}-\langle p_{T}\rangle_{\alpha}^{2},$ (10)
where the averages are taken over all particles from all events belonging to the given $N_{\mathrm{part}}$ bin.
@@
To investigate the response of transverse-momentum fluctuations for different particle species, the variance of the single-particle transverse-momentum distribution is evaluated separately for each species $i$ within a fixed $N_{\mathrm{part}}$ class,
$\delta p_{T,i}^{2}=\Big\langle\!\Big\langle\langle p_{T}^{2}\rangle_{i}-\langle p_{T}\rangle_{i}^{2}\Big\rangle\!\Big\rangle.$ (12)
Here, $\langle\cdots\rangle_{i}$ denotes an average over all particles of species $i$ and over all events belonging to the given $N_{\mathrm{part}}$ bin. The index $i$ labels the particle species (e.g. $\pi$ or $p$).
@@
Figure 16: Ensemble-averaged single-particle transverse-momentum variance $\delta p_{T,i}^{2}$ as a function of $N_{\mathrm{part}}$ for different particle species in C+C collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
Figure 17: Ensemble-averaged single-particle transverse-momentum variance $\delta p_{T,i}^{2}$ as a function of $N_{\mathrm{part}}$ for different particle species in C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
@@
To characterize the magnitude and event-by-event fluctuations of the flow harmonics, we evaluate moments of the event-wise flow coefficients using the $Q$-vector formalism. For each event, the quantity $v_{n}^{2}=|Q_{n}|^{2}/M^{2}$ is constructed, where $Q_{n}=\sum_{i}e^{in\phi_{i}}$ and $M$ denotes the multiplicity of selected particles. The typical magnitude of the flow harmonics is characterized by the root-mean-square value $v_{n}\{2\}=\sqrt{\langle v_{n}^{2}\rangle}$, while the fluctuation strength of a given harmonic is quantified by the variance $\langle v_{n}^{4}\rangle-\langle v_{n}^{2}\rangle^{2}$, with angle brackets denoting an average over events at fixed $N_{\mathrm{part}}$. The quantity $v_{n}\{2\}$ corresponds to the root-mean-square flow coefficient and is sensitive to both the mean flow and its event-by-event fluctuations. Only events with charged-particle multiplicity $N_{\mathrm{ch}}(|\eta|<2)>4$ are included in the analysis.
@@
Figure 18: Root-mean-square flow magnitude $v_{1}\{2\}=\sqrt{\langle v_{1}^{2}\rangle}$ and the corresponding variance $\langle v_{1}^{4}\rangle-\langle v_{1}^{2}\rangle^{2}$ as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
@@
Figure 19: Root-mean-square flow magnitude $v_{1}\{2\}=\sqrt{\langle v_{2}^{2}\rangle}$ and the corresponding variance $\langle v_{2}^{4}\rangle-\langle v_{2}^{2}\rangle^{2}$ as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
@@
Figure 20: Root-mean-square flow magnitude $v_{3}\{2\}=\sqrt{\langle v_{3}^{2}\rangle}$ and the corresponding variance $\langle v_{3}^{4}\rangle-\langle v_{3}^{2}\rangle^{2}$ as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
@@
Figures 18, 19, and 20 present the root-mean-square (RMS) flow magnitude and the corresponding fluctuation strength for the first-, second-, and third-order flow harmonics, evaluated using ensemble-averaged moments at fixed $N_{\mathrm{part}}$. We observe that $\alpha$ clustering leads to a significant enhancement of the RMS flow magnitude, particularly at large $N_{\mathrm{part}}$ ($\gtrsim 70$). In contrast, the variance $\langle v_{n}^{4}\rangle-\langle v_{n}^{2}\rangle^{2}$ of individual harmonics remains small and is not clearly distinguishable within the current statistical precision. This indicates that $\alpha$ clustering primarily modifies the average initial geometry and its collective response, while inducing only modest changes in the magnitude of single-harmonic flow fluctuations. Consequently, precise experimental measurements of flow observables in this regime could provide valuable constraints on the presence and strength of $\alpha$ clustering in light nuclei.
@@
Finally, ensemble-averaged symmetric cumulants of the flow harmonics are evaluated. Figure 21 shows $\mathrm{SC}(v_{1}^{2},v_{3}^{2})$ and $\mathrm{SC}(v_{2}^{2},v_{3}^{2})$ for C+Pb collisions, where the cumulants are constructed from ensemble-averaged moments over all particles and events at fixed $N_{\mathrm{part}}$. Both configurations yield a positive $\mathrm{SC}(v_{1}^{2},v_{3}^{2})$, with no statistically significant distinction between them. The $\mathrm{SC}(v_{2}^{2},v_{3}^{2})$ values are consistent with zero for both the $\alpha$-clustering and Woods–Saxon cases.